A rock is thrown horizontally from the top of a cliff 88 m high, with a horizontal speed

of 25 m/s.With what velocity does the rock hit?


1
Expert's answer
2022-05-19T05:43:17-0400

"v_x=25~\frac ms,"

"v_y=\sqrt{2gh}=41.5~\frac ms,"

"v=\sqrt{v_x^2+v_y^2}=48.5~\frac ms."